OpenStudy (anonymous):

A dehumidifier with a resistance of 16Ω draws 5.4A when connected to an electrical outlet. Find the power consumed by the appliance.

OpenStudy (anonymous):

Power = P\[P = I^2 R\]

OpenStudy (anonymous):

1382?

OpenStudy (anonymous):

Power = (5.4)^2 * 16 = 466.56 Watts

OpenStudy (anonymous):

Oh. I did that backwards. Thanks